The equilibrium constant K for the reaction below is is 7.755x105. CODE 4th attempt Part 1 (1 point) W See Periodic Table TheI got the first one correct but not the last 2.
Answer #1

C initially D + E [c] = [0] = [E] 1•7110 x 10-3 M. KES Oc = [DJ [E] - (1•7110x10-3)2 H [c] - Clo7H0 X10-3) -1 E 107110 X 10= [DJ TE] [C] (1:7/10 x 10-3-212 (107110X10-3+2) L OL 01 *0-3 70755 x 10-5 = x2 - 3,4220x + 2.927521 X 10 (107110X10-3+2) SolHope you will like my answer.